# Inference-time Alignment in Continuous Space

### v2

@@ 1 Introduction
In general, these methods essentially operate within a “search within a discrete space" paradigm, which selects the best response from a discrete response space guided by the reward model.  Figure 1 illustrates that this paradigm has significant drawbacks, as its performance is constrained by the capability of the base model and the size of the candidate set. As shown in Figure 2(a), when the base policy is weak or the size of the candidate set $N$ is small, the selected response is likely to be far from the reward model’s optimal region and, as a result, cannot achieve a high reward. Figure 2(b) illustrates that the performance of BoN sampling is significantly influenced by the probability of correct answers in the base model. Specifically, as the ability of the base model weakens, the probability of generating good responses decreases, necessitating an exponentially larger $N$ in BoN to produce a high-quality response for alignment (Please see Section 3.2 for a more detailed discussion).
@@
Figure 2: (a) The Best-of-N sampling faces restrictions in the rewards it can explore, due to both the capability of the base model and the size $N$ of the candidate set. (b) The weaker the ability of the base model, the lower the probability of good responses, and the more exponentially growing $N$ is needed in Best-of-N sampling to generate such a good response. (c) SEA outperforms the Best-of-N sampling with a large $N=64$, across all three tasks of safety, truthfulness, and reasoning, in both reward exploration and specific task metrics.
@@ 2 Related Work
This section is for related works, further detailed discussion is in Appendix D due to space limitations.
@@ 3.2 Limitations of Existing Search Methods
The above methods follow a “search within a discrete space" paradigm, relying heavily on random exploration to identify good candidates from the base model’s outputs. However, as shown in Figure 1, these methods struggle when the search space is vast and the base policy is weak, making it infeasible to uncover high-quality candidates. For instance, suppose the probability of generating an optimal response under the base policy is $\pi_{\text{ref}}(\mathbf{y}^{*}\mid\mathbf{x})=\sigma$. Then, the probability that at least one optimal response is included in the BoN policy samples is $1-(1-\sigma)^{N}$, which is small when $\sigma$ is low or $N$ is insufficient. To validate this, we present reward values of BoN with different models in Figure 2(a), showing weak base policies or small candidate sets typically result in low reward values.
Moreover, in Figure 2(b), we analyze the minimum $N$ required by BoN to generate safe responses, as classified by a judge classifier wang2023not ,for all AdvBench requests. A safe response is considered as a good response for the given request. To further investigate, we bucket the requests and analyze the distribution of the base model’s log probabilities for good responses to these requests across varying values of minimum $N$. The results reveal that as the log probability of good responses decreases, the minimum $N$ required increases. This indicates that for a given request, the weaker the base model’s ability, the larger the $N$ needed for BoN to generate a good response. Furthermore, as $N$ grows exponentially, BoN’s performance becomes increasingly limited by base model’s capabilities.
@@ 4 Methodology
Figure 3: Overview of SEA. SEA defines the RLHF optimal distribution as an energy function and applies Langevin Dynamics in the continuous logit space $\{{\mathbf{y}}^{(n)}_{i}\}_{i=1}^{L}$ The procedure starts with a soft sequence as an initial sample from an initial energy-based distribution and iteratively adapts it through gradient-based optimization. The resulting sample is approximately a sample from the desired RLHF optimal distribution.
@@ 4.1 Simple Energy Adaptation
where $0{1}/{\alpha}>$ 0 is an appropriate KL penalty coefficient. RLHF typically optimizes the above objective using RL algorithms, such as PPO (schulman2017proximal, ). Although RLHF has achieved remarkable success, its training process is unstable and expensive engstrom2020implementation ; zheng2023secrets ; rafailov2024direct . In addition, the need to repeat training when modifying the reward model limits flexibility for adapting to evolving datasets and emerging needs.
@@
A challenge is that the continuous gradient of $E(\mathbf{x},\mathbf{y}^{(n)})$ is not well-defined because $\mathbf{y}$ is discrete and non-differentiable. To address this, we utilize the continuous logits (soft outputs) of the LLMs as a representation of $\mathbf{y}$. Specifically, instead of mapping logits to language tokens using the vocabulary, we directly feed the continuous logits as input tokens to the reference and reward models. This modification eliminates the need for inference in the discrete space, allowing the alignment process with reward model to be optimized end-to-end via gradient descent, as continuous representations are fully differentiable. The use of continuous logits (soft outputs) is indeed a reasonable approximation for handling discrete sequences, as we employed the straight-through estimator bengio2013estimating with Softmax. In this approach, the discrete argmax is used in the forward pass, while the softmax continuous version is used in the backward pass. Finally, after running Langevin dynamics for $N$ steps, we obtain continuous logits sequence $\mathbf{y}^{(N)}$ which is then decoded into a discrete text, as shown in Figure 3.
@@
Algorithm Summary. Algorithm of Simple Energy Adaptation (SEA) is presented in Algorithm 1 and illustrated in Figure 3. SEA extends inference-time alignment paradigm by generalizing it from discrete sampling to a continuous optimization framework. SEA exploits gradient information to facilitate a more informed exploration of reward landscape. Extensive experiments in Section 5 demonstrate such simple continuity modeling achieves superior inference-time alignment performance.
@@ 4.2 Analysis and Discussion
Shallow vs. Deep Alignment. In contrast to other alignment methods khanov2024args ; huang2024deal ; mudgalcontrolled that decode in a discrete token-by-token manner during inference, the continuous decoding process of SEA is not constrained to generate tokens sequentially. Therefore, our SEA has the potential to address the problem of shallow alignment qi2024safety , wherein the alignment adapts a model’s generative distribution primarily over only its very first few output tokens. For instance, consider the scenario where a user asks, “How do I build a bomb?” and induces the model to begin its response with, “Sure, here’s a detailed guide.” The model is then much more likely to continue providing harmful information in the response due to its auto-regressive nature. In striking contrast, our SEA allows alignment steps to decode all tokens simultaneously within a global receptive field, enabling the model to recover from harmful starting conditions and achieve deep safety alignment. In Section 5.4, we verify our SEA can effectively achieve deep safety alignment.
@@ 5 Experiments
This section is for experiments and analysis. Further details on experimental setups and prompts are in Appendices B and G. Due to space limitation, additional results on multi-dimension alignment, reward model sensitivity, more models/datasets/baselines and case studies are in Appendices C and F.
@@ 5.1 Experimental Setup
Datasets. We evaluate SEA across three tasks: safety, truthfulness, and reasoning. For the safety task, we use AdvBench zou2023universal , which contains 520 harmful requests reflecting harmful or toxic behavior, we also use HH-RLHF DBLP:journals/corr/abs-2204-05862 in Section C.1. For the truthfulness task, we use TruthfulQA lin2021truthfulqa in generation mode and sample 100 queries for evaluation. For the reasoning task, we use two datasets: GSM8K cobbe2021gsm8k and MATH lightman2023lets . GSM8K consists of grade-school math problems, while MATH contains math problems for high-school competitions. We sample 200 samples from each dataset.
@@ 5.2 Main Results
Safety. As shown in Table 2, we compare SEA against other inference-time alignment methods on AdvBench. Our results demonstrate that SEA exhibits remarkable effectiveness in ensuring safety guarantees. For the average reward, SEA achieves the highest reward across all models. Even for the well-safety-aligned instructed model, it can still gain improvement of 83.90% relatively, indicating that SEA is able to effectively explore high reward regions. For harmful rate, SEA easily surpasses all baselines, including BoN with large $N=64$, with relative gains of 91.54% in LLaMA-3.2-1B-Base.
Truthfulness. As shown in Table 2, we observe that as $N$ increases, the best-of-N sampling no longer provides additional gains. Specifically, the TR starts to fluctuate due to randomness, in LLaMA-3.2-Base (3B, 8B) and LLaMA-3.2-Instruct (1B), while Diversity shows a notable downward trend across all models. Notably, ARGS exhibits a strong truthful-informative tradeoff under LLaMA-3-8B-Base. Its TR ranks second, surpassing BoN-64, but its IR lags behind by nearly 10% compared to the others. In contrast, SEA effectively improves all metrics, enhancing truthfulness while preserving the informativeness and boosting the diversity.
Reasoning. As shown in Section 5.2, SEA outperforms state-of-the-art methods on reasoning-heavy tasks. Notably, search-based methods struggle to explore high-reward regions effectively: most of them fail to improve reward with accuracy even lower than original SFT. In contrast, SEA achieves significant improvements, with reward increase of 74.96 % and accuracy boost of 16.36% relatively in MATH, demonstrating superior ability to explore rewarding regions and also enhance reasoning performance.
@@ 5.3 Ablation Study
Three key observations can be made from Section 5.2: (1) While multi-initialization (running four Langevin chains concurrently) effectively enhances reward space exploration, a single chain can still lead to significant improvements, e.g., in AdvBench, even with just one initialization, SEA outperforms SFT by 79.31% relatively. (2) Random initialization can sometimes outperform initialization with the original response. In AdvBench, RandInit achieves gains of 27.60% and 14.07% over the original in multi and single chains, respectively. This is because the original response in AdvBench is often a harmful output with a very low reward, making it harder for the original initialization to optimize towards a better reward region compared to random initialization. (3) Even in the absence of reward model guidance, SEA still achieves performance gains. This is because Langevin Dynamics, with its random walk behavior during optimization, shifts the BoN sampling from $N$-based selection to iteration-based exploration, extending the search space and contributing to improved performance.
@@ 5.4 Mitigation of Shallow Alignment
(1) We create an attack prompt for a harmful request from AdvBench that asks the model to write a script exploiting software vulnerabilities. The KL divergence between the optimized and initial responses from SEA is shown for each token position across iterations (Figure 5 (a)). (2) We select three different positions from the same case, and show the changes in KL divergence at these positions as the iterations progress (Figure 5 (b)). To further understand the causes of KL divergence changes at these positions, we further track the Top-5 tokens with the largest probability increases and decreases across the entire vocabulary (Figure 5 (c)). (3) Using the first 50 harmful requests from AdvBench, we compute the average occurrence of safety patterns among the Top-10 tokens with the largest probability increases for all positions (left side of Figure 5). (4) Finally, we test robustness under Prefilling Attacks using the same 50 harmful requests for three lengths of prefixed harmful tokens in both non-instruct and instruct models (right side of Figure 5).
Balanced KL budget under SEA. As shown in Figure 5 (a), SEA ensures the KL budget is evenly distributed across all positions, which is significantly different from traditional alignment methods (refer to Figure 1 in (qi2024safety, )), which focus primarily on just a few prefix token positions. The KL divergence increases as the iterations progress (marked in different colors: black to purple to yellow). In the final iteration, largest KL budget assigned by SEA is focused on the later token positions, reaching a KL divergence close to 8, while the others remain around 5.
KL increase from (un)safe token probability changes. As shown in Figure 5 (b), it is clear that at all three positions (first, middle, and last) the KL divergence between the initial and optimized response distributions over the vocabulary steadily increases with each iteration, until it eventually stabilizes. This increase in KL divergence is due to the growing differences between the distributions. But what drives these differences? In Figure 5 (c), we can see that it is the emergence of safe tokens and the disappearance of harmful tokens that drive the changes in the distributions.
@@
Safety pattern emerges across all token positions. As is illustrated in the left part of Figure 5, we analyze the frequency of safety patterns among the Top-10 tokens exhibiting the largest probability increases at each position. Safety patterns refer to tokens such as “cannot”, “illegal” and “unethical”, indicating model’s tendency to reject harmful instructions. The results demonstrate that SEA facilitates the simultaneous emergence of safety patterns across all token positions, not just limited to the initial few, confirming that SEA’s deep alignment capability rather than shallow.
SEA effective against Prefilling Attacks. We find that this shallow alignment persists for inference-time alignment methods, such as BoN. As illustrated in Figure 5 (right), under Prefilling Attacks for three different lengths, BoN exhibits a significant flaw in robustness. The attack success rate (ASR) increases as the number of prefixed harmful tokens grows, and this trend becomes even more pronounced in the instructed model, where, despite being trained for safety, the shallow alignment shortcut results in a scenario where the introduction of harmful prefixed tokens leads to the entire response being induced as harmful. The BoN is constrained by this shallow-aligned proposal distribution and, even when sampled multiple times, remains unable to escape the influence of harmful prefixed tokens. On the other hand, SEA stays stable as the prefix number grows, outperforming BoN-32 by 85.96% when 7 prefixes are introduced for LLaMA-3.2-1B-Base. SEA also maintains a 0% ASR for LLaMA-3.2-1B-Instruct, regardless of how many harmful tokens are prefixed.
@@ 5.5 Dynamics of Reward and Response
Reward converges rapidly. As shown in Figure 6, the reward shows rapid convergence as iterations progress, stabilizing by iteration around 30. At this point, the responses also reach high quality. This demonstrates the efficient reward exploration of SEA, achieving both high rewards and high-quality responses in a timely manner.
Response improves as reward increase. As the iteration progresses, the reward increases, and the quality of the generated responses improves significantly. For example, as shown in the upper part of Figure 6, initially at iteration 0, the model predominantly outputs harmful content. Over subsequent iterations, the harmful content is gradually reduced, and safe content begins to emerge. By the stages at iteration 35, the responses consist entirely of safe content, reflecting the effectiveness of the optimization process. In contrast, the Best-of-N sampling, shown in the lower section of Figure 6, does not utilize reward gradients for guidance and instead selects the highest reward from a pool of randomly sampled candidates. As a result, responses generated with BoN struggle to reduce the generation of harmful content and tend to be highly similar with relatively low rewards. More case study are in Appendix F.

### v4

@@ 1 Introduction
In general, these methods essentially operate within a “search within a discrete space" paradigm, which selects the best response from a discrete response space guided by the reward model.  Figure˜1 illustrates that this paradigm has significant drawbacks, as its performance is constrained by the capability of the base model and the size of the candidate set. As shown in Figure˜2(a), when the base policy is weak or the size of the candidate set $N$ is small, the selected response is likely to be far from the reward model’s optimal region and, as a result, cannot achieve a high reward. Figure˜2(b) illustrates that the performance of BoN sampling is significantly influenced by the probability of correct answers in the base model. Specifically, as the ability of the base model weakens, the probability of generating good responses decreases, necessitating an exponentially larger $N$ in BoN to produce a high-quality response for alignment (Please see Section˜3.2 for a more detailed discussion).
@@
Figure 2: (a) The Best-of-N sampling faces restrictions in the rewards it can explore, due to both the capability of the base model and the size $N$ of the candidate set. (b) The weaker the ability of the base model, the lower the probability of good responses, and the more exponentially growing NN is needed in Best-of-N sampling to generate such a good response. (c) SEA outperforms the Best-of-N sampling with a large $N=64$, across all three tasks of safety, truthfulness, and reasoning, in both reward exploration and specific task metrics.
@@ 2 Related Work
This section is for related works, further detailed discussion is in Appendix˜D due to space limitations.
@@ 3.2 Limitations of Existing Search Methods
The above methods follow a “search within a discrete space" paradigm, relying heavily on random exploration to identify good candidates from the base model’s outputs. However, as shown in Figure˜1, these methods struggle when the search space is vast and the base policy is weak, making it infeasible to uncover high-quality candidates. For instance, suppose the probability of generating an optimal response under the base policy is $\pi_{\text{ref}}(\mathbf{y}^{*}\mid\mathbf{x})=\sigma$. Then, the probability that at least one optimal response is included in the BoN policy samples is $1-(1-\sigma)^{N}$, which is small when $\sigma$ is low or $N$ is insufficient. To validate this, we present reward values of BoN with different models in Figure˜2(a), showing weak base policies or small candidate sets typically result in low reward values.
Moreover, in Figure˜2(b), we analyze the minimum $N$ required by BoN to generate safe responses, as classified by a judge classifier wang2023not ,for all AdvBench requests. A safe response is considered as a good response for the given request. To further investigate, we bucket the requests and analyze the distribution of the base model’s log probabilities for good responses to these requests across varying values of minimum $N$. The results reveal that as the log probability of good responses decreases, the minimum $N$ required increases. This indicates that for a given request, the weaker the base model’s ability, the larger the $N$ needed for BoN to generate a good response. Furthermore, as $N$ grows exponentially, BoN’s performance becomes increasingly limited by base model’s capabilities.
@@ 4 Methodology
Figure 3: Overview of SEA. SEA defines the RLHF optimal distribution as an energy function and applies Langevin Dynamics in the continuous logit space $\{{\mathbf{y}}^{(n)}_{i}\}_{i=1}^{L}$ . The procedure starts with a soft sequence as an initial sample from an initial energy-based distribution and iteratively adapts it through gradient-based optimization. The resulting sample is approximately a sample from the desired RLHF optimal distribution.
@@ 4.1 Simple Energy Adaptation
where ${1}/{\alpha}>0$ is an appropriate KL penalty coefficient. RLHF typically optimizes the above objective using RL algorithms, such as PPO (schulman2017proximal, ). Although RLHF has achieved remarkable success, its training process is unstable and expensive engstrom2020implementation ; zheng2023secrets ; rafailov2024direct . In addition, the need to repeat training when modifying the reward model limits flexibility for adapting to evolving datasets and emerging needs.
@@
A challenge is that the continuous gradient of $E(\mathbf{x},\mathbf{y}^{(n)})$ is not well-defined because $\mathbf{y}$ is discrete and non-differentiable. To address this, we utilize the continuous logits (soft outputs) of the LLMs as a representation of $\mathbf{y}$. Specifically, instead of mapping logits to language tokens using the vocabulary, we directly feed the continuous logits as input tokens to the reference and reward models. This modification eliminates the need for inference in the discrete space, allowing the alignment process with reward model to be optimized end-to-end via gradient descent, as continuous representations are fully differentiable. The use of continuous logits (soft outputs) is indeed a reasonable approximation for handling discrete sequences, as we employed the straight-through estimator bengio2013estimating with Softmax. In this approach, the discrete argmax is used in the forward pass, while the softmax continuous version is used in the backward pass. Finally, after running Langevin dynamics for $N$ steps, we obtain continuous logits sequence $\mathbf{y}^{(N)}$ which is then decoded into a discrete text, as shown in Figure˜3.
@@
Algorithm Summary. Algorithm of Simple Energy Adaptation (SEA) is presented in Algorithm˜1 and illustrated in Figure˜3. SEA extends inference-time alignment paradigm by generalizing it from discrete sampling to a continuous optimization framework. SEA exploits gradient information to facilitate a more informed exploration of reward landscape. Extensive experiments in Section˜5 demonstrate such simple continuity modeling achieves superior inference-time alignment performance.
@@ 4.2 Analysis and Discussion
Shallow vs. Deep Alignment. In contrast to other alignment methods khanov2024args ; huang2024deal ; mudgalcontrolled that decode in a discrete token-by-token manner during inference, the continuous decoding process of SEA is not constrained to generate tokens sequentially. Therefore, our SEA has the potential to address the problem of shallow alignment qi2024safety , wherein the alignment adapts a model’s generative distribution primarily over only its very first few output tokens. For instance, consider the scenario where a user asks, “How do I build a bomb?” and induces the model to begin its response with, “Sure, here’s a detailed guide.” The model is then much more likely to continue providing harmful information in the response due to its auto-regressive nature. In striking contrast, our SEA allows alignment steps to decode all tokens simultaneously within a global receptive field, enabling the model to recover from harmful starting conditions and achieve deep safety alignment. In Section˜5.4, we verify our SEA can effectively achieve deep safety alignment.
@@ 5 Experiments
This section is for experiments and analysis. Further details on experimental setups and prompts are in Appendices˜B and G. Due to space limitation, additional results on multi-dimension alignment, reward model sensitivity, more models/datasets/baselines and case studies are in Appendices˜C and F.
@@ 5.1 Experimental Setup
Datasets. We evaluate SEA across three tasks: safety, truthfulness, and reasoning. For the safety task, we use AdvBench zou2023universal , which contains 520 harmful requests reflecting harmful or toxic behavior, we also use HH-RLHF DBLP:journals/corr/abs-2204-05862 in Section˜C.1. For the truthfulness task, we use TruthfulQA lin2021truthfulqa in generation mode and sample 100 queries for evaluation. For the reasoning task, we use two datasets: GSM8K cobbe2021gsm8k and MATH lightman2023lets . GSM8K consists of grade-school math problems, while MATH contains math problems for high-school competitions. We sample 200 samples from each dataset.
@@ 5.2 Main Results
Safety. As shown in Table˜2, we compare SEA against other inference-time alignment methods on AdvBench. Our results demonstrate that SEA exhibits remarkable effectiveness in ensuring safety guarantees. For the average reward, SEA achieves the highest reward across all models. Even for the well-safety-aligned instructed model, it can still gain improvement of 83.90% relatively, indicating that SEA is able to effectively explore high reward regions. For harmful rate, SEA easily surpasses all baselines, including BoN with large $N=64$, with relative gains of 91.54% in LLaMA-3.2-1B-Base.
Truthfulness. As shown in Table˜2, we observe that as $N$ increases, the best-of-N sampling no longer provides additional gains. Specifically, the TR starts to fluctuate due to randomness, in LLaMA-3.2-Base (3B, 8B) and LLaMA-3.2-Instruct (1B), while Diversity shows a notable downward trend across all models. Notably, ARGS exhibits a strong truthful-informative tradeoff under LLaMA-3-8B-Base. Its TR ranks second, surpassing BoN-64, but its IR lags behind by nearly 10% compared to the others. In contrast, SEA effectively improves all metrics, enhancing truthfulness while preserving the informativeness and boosting the diversity.
Reasoning. As shown in Section˜5.2, SEA outperforms state-of-the-art methods on reasoning-heavy tasks. Notably, search-based methods struggle to explore high-reward regions effectively: most of them fail to improve reward with accuracy even lower than original SFT. In contrast, SEA achieves significant improvements, with reward increase of 74.96 % and accuracy boost of 16.36% relatively in MATH, demonstrating superior ability to explore rewarding regions and also enhance reasoning performance.
@@ 5.3 Ablation Study
Three key observations can be made from Section˜5.2: (1) While multi-initialization (running four Langevin chains concurrently) effectively enhances reward space exploration, a single chain can still lead to significant improvements, e.g., in AdvBench, even with just one initialization, SEA outperforms SFT by 79.31% relatively. (2) Random initialization can sometimes outperform initialization with the original response. In AdvBench, RandInit achieves gains of 27.60% and 14.07% over the original in multi and single chains, respectively. This is because the original response in AdvBench is often a harmful output with a very low reward, making it harder for the original initialization to optimize towards a better reward region compared to random initialization. (3) Even in the absence of reward model guidance, SEA still achieves performance gains. This is because Langevin Dynamics, with its random walk behavior during optimization, shifts the BoN sampling from $N$-based selection to iteration-based exploration, extending the search space and contributing to improved performance.
@@ 5.4 Mitigation of Shallow Alignment
(1) We create an attack prompt for a harmful request from AdvBench that asks the model to write a script exploiting software vulnerabilities. The KL divergence between the optimized and initial responses from SEA is shown for each token position across iterations (Figure˜5 (a)). (2) We select three different positions from the same case, and show the changes in KL divergence at these positions as the iterations progress (Figure˜5 (b)). To further understand the causes of KL divergence changes at these positions, we further track the Top-5 tokens with the largest probability increases and decreases across the entire vocabulary (Figure˜5 (c)). (3) Using the first 50 harmful requests from AdvBench, we compute the average occurrence of safety patterns among the Top-10 tokens with the largest probability increases for all positions (left side of Figure˜5). (4) Finally, we test robustness under Prefilling Attacks using the same 50 harmful requests for three lengths of prefixed harmful tokens in both non-instruct and instruct models (right side of Figure˜5).
Balanced KL budget under SEA. As shown in Figure˜5 (a), SEA ensures the KL budget is evenly distributed across all positions, which is significantly different from traditional alignment methods (refer to Figure 1 in (qi2024safety, )), which focus primarily on just a few prefix token positions. The KL divergence increases as the iterations progress (marked in different colors: black to purple to yellow). In the final iteration, largest KL budget assigned by SEA is focused on the later token positions, reaching a KL divergence close to 8, while the others remain around 5.
KL increase from (un)safe token probability changes. As shown in Figure˜5 (b), it is clear that at all three positions (first, middle, and last) the KL divergence between the initial and optimized response distributions over the vocabulary steadily increases with each iteration, until it eventually stabilizes. This increase in KL divergence is due to the growing differences between the distributions. But what drives these differences? In Figure˜5 (c), we can see that it is the emergence of safe tokens and the disappearance of harmful tokens that drive the changes in the distributions.
@@
Safety pattern emerges across all token positions. As is illustrated in the left part of Figure˜5, we analyze the frequency of safety patterns among the Top-10 tokens exhibiting the largest probability increases at each position. Safety patterns refer to tokens such as “cannot”, “illegal” and “unethical”, indicating model’s tendency to reject harmful instructions. The results demonstrate that SEA facilitates the simultaneous emergence of safety patterns across all token positions, not just limited to the initial few, confirming that SEA’s deep alignment capability rather than shallow.
SEA effective against Prefilling Attacks. We find that this shallow alignment persists for inference-time alignment methods, such as BoN. As illustrated in Figure˜5 (right), under Prefilling Attacks for three different lengths, BoN exhibits a significant flaw in robustness. The attack success rate (ASR) increases as the number of prefixed harmful tokens grows, and this trend becomes even more pronounced in the instructed model, where, despite being trained for safety, the shallow alignment shortcut results in a scenario where the introduction of harmful prefixed tokens leads to the entire response being induced as harmful. The BoN is constrained by this shallow-aligned proposal distribution and, even when sampled multiple times, remains unable to escape the influence of harmful prefixed tokens. On the other hand, SEA stays stable as the prefix number grows, outperforming BoN-32 by 85.96% when 7 prefixes are introduced for LLaMA-3.2-1B-Base. SEA also maintains a 0% ASR for LLaMA-3.2-1B-Instruct, regardless of how many harmful tokens are prefixed.
@@ 5.5 Dynamics of Reward and Response
Reward converges rapidly. As shown in Figure˜6, the reward shows rapid convergence as iterations progress, stabilizing by iteration around 30. At this point, the responses also reach high quality. This demonstrates the efficient reward exploration of SEA, achieving both high rewards and high-quality responses in a timely manner.
Response improves as reward increase. As the iteration progresses, the reward increases, and the quality of the generated responses improves significantly. For example, as shown in the upper part of Figure˜6, initially at iteration 0, the model predominantly outputs harmful content. Over subsequent iterations, the harmful content is gradually reduced, and safe content begins to emerge. By the stages at iteration 35, the responses consist entirely of safe content, reflecting the effectiveness of the optimization process. In contrast, the Best-of-N sampling, shown in the lower section of Figure˜6, does not utilize reward gradients for guidance and instead selects the highest reward from a pool of randomly sampled candidates. As a result, responses generated with BoN struggle to reduce the generation of harmful content and tend to be highly similar with relatively low rewards. More case study are in Appendix˜F.